Session: May 2004 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K. Borkowski at Mon Jun 7 05:58:39 2004) Success :
No known problems.

Medicina (scalambra at Mon Jun 7 06:02:01 2004) Minor failures :

Onsala (Michael Lindqvist at Mon Jun 7 06:15:21 2004) Success :
No known problems apart from some high parity errors.

Westerbork (Tony Foley at Mon Jun 7 08:10:31 2004) Minor failures : PARI HA-E
Wb used in phased array mode. NOTE: we re-cabled from Mk5A to Mk4 for this. Parity errors slightly high. Scans from 22:28 to 22:38 out of Hour Angle range.

Jodrell_Bank (Alastair Gunn at Mon Jun 7 10:00:27 2004) Success :
MkII was used. No known problems.

Effelsberg (A. Kraus at Mon Jun 7 11:16:49 2004) Success :

Noto (P. Cassaro at Tue Jun 8 09:55:01 2004) Success :
The tape recorder in Noto is still broken. The experiment has been recorded with MK5.

Robledo (Cristina Garcia Miro at Fri Jun 11 14:25:10 2004) Minor failures : RECO LATE RFI
VC01 completely degraded by L-band RFI (1660-1670 GHz). First 4 sources lost, 5th partially lost. Tape positioning incorrect for all passes on first tape (+1000 feet). Not enough time provided in schedule for tape change. As a result first source of second tape not recorded and following sources recorded 600 feet off.